how many different ways can 5 people be seated on a bench if 2 of the people must sit next to eachother?
Answer by Edwin McCravy(20054)

how many different ways can 5 people be seated on a bench if 2 of the people must sit next to eachother?
Think of handcuffing the two together that mist sit next to each
other and then you only have 4 "things" to seat, 3 single people
and 1 couple handcuffed together.

There are 2 ways to handcuff them together.
For each of the two ways to handcuff them together, there are
4! ways to seat the four "things", so the answer is 

2*4! = 2*4*3*2*1 = 48 ways.

Edwin